# From Prompt to Pipeline: Large Language Models for Scientific Workflow Development in Bioinformatics

### v1

@@ 4.1.2 Workflow Development using LLMs for Nextflow:
We finally give the same prompt to DeepSeek-V3. It also captures both placement-only and search-plus-placement modes and structures the pipeline into logically ordered modules. It specifies necessary input parameters, including reference alignment, rooted phylogenetic tree, evolutionary model, taxonomy file, and HMM profiles, and introduces validation steps to ensure input integrity. The outlined modules, covering HMM-based sequence filtering (hmmsearch), alignment (MAFFT or HMMER), placement (EPA-NG or pplacer), and classification (gappa or taxit), are consistent with tools used in the official pipeline. DeepSeek-V3 additionally incorporates optional visualization using tools such as ETE3, R, and Python, and provides a pseudocode-level implementation using Nextflow syntax that demonstrates channel logic and parameter control.
All three LLMs, successfully reconstruct a modular Nextflow workflow for phylogenetic placement, capturing both operational modes and correctly identifying core steps such as input validation, sequence alignment, phylogenetic placement, classification, and reporting. GPT-4o provides a technically accurate and semantically faithful outline closely aligned with the nf-core implementation. Gemini 2.5 Flash presents a well-organized modular architecture using intuitive labels and standard tools, though it omits some nf-core-specific features like execution metadata. DeepSeek-V3 stands out by offering the most implementation-ready response, including detailed parameter definitions, alternative tools (e.g., pplacer, taxit), and a fully structured Nextflow pseudocode, making it especially valuable for advanced users or developers seeking practical scaffolding. While all three are valid, DeepSeek-V3 provides the most comprehensive and adaptable solution, GPT-4o ensures the highest fidelity to the reference workflow, and Gemini excels in clarity for novices. Therefore, DeepSeek-V3 is best suited for expert-driven implementation, GPT-4o for faithful documentation reproduction, and Gemini for accessible modular reasoning.
@@ 8 Conclusion
[…]
Supplementary information All supplementary files can be obtained using [88]
[…]
Acknowledgements This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).

### v2

@@ 4.1.2 Workflow Development using LLMs for Nextflow:
We finally give the same prompt to DeepSeek-V3. It also captures both placement-only and search-plus-placement modes and structures the pipeline into logically ordered modules. It specifies necessary input parameters, including reference alignment, rooted phylogenetic tree, evolutionary model, taxonomy file, and HMM profiles, and introduces validation steps to ensure input integrity. The outlined modules, covering HMM-based sequence filtering (hmmsearch), alignment (MAFFT or HMMER), placement (EPA-NG or pplacer), and classification (gappa or taxit), are consistent with tools used in the official pipeline. DeepSeek-V3 additionally incorporates optional visualization using tools such as ETE3, R, and Python, and provides a pseudocode-level implementation using Nextflow syntax.
All three LLMs, successfully reconstruct a modular Nextflow workflow for phylogenetic placement, capturing both operational modes and correctly identifying core steps such as input validation, sequence alignment, phylogenetic placement, classification, and reporting. GPT-4o provides a technically accurate and semantically faithful outline closely aligned with the nf-core implementation. Gemini 2.5 Flash presents a well-organized modular architecture using intuitive labels and standard tools, though it omits some nf-core-specific features like execution metadata. DeepSeek-V3 stands out by offering the most implementation-ready response, including detailed parameter definitions, alternative tools (e.g., pplacer, taxit), and a fully structured Nextflow pseudocode, making it especially valuable for advanced users or developers seeking practical scaffolding. While all three are valid, DeepSeek-V3 provides the most comprehensive and adaptable solution, GPT-4o ensures the highest fidelity to the reference workflow, and Gemini excels in clarity for novices. Therefore, DeepSeek-V3 is best suited for expert-driven implementation, GPT-4o for documentation reproduction, and Gemini for accessible modular reasoning.
@@ 8 Conclusion
Supplementary information
All supplementary files can be obtained using [93]
Acknowledgment
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).
Data availability statement
All supplementary files can be obtained using [93]